# Human Perception-Guided Pretraining for Increased Generalization

### v2

@@ 5 Results
       ResNet152 0.852$\pm$0.001 0.864$\pm$0.002 0.864$\pm$0.002
       Inception-V4 0.848$\pm$0.002 0.846$\pm$0.003 0.856$\pm$0.001
       EfficientNet-7b 0.827$\pm$0.002 0.827$\pm$0.002 0.845$\pm$0.005
5.1 Answering RQ1 (improvement of generalization)
MENTOR pre-training compares very favorably to models initialized with ImageNet weights, as well as to models trained with a state-of-the-art human-perception training paradigm (CYBORG) using the same data and the same human saliency maps. In six out of nine domain-architecture combinations, MENTOR achieves statistically significantly better average AUROC (across 10 independent training runs) than CYBORG- and pure cross-entropy-based training (cf. Figs 3-5).
Looking only at top-three average AUROC scores (Tab. 1), MENTOR pre-training outperforms both models initialized with ImageNet-sourced weights and (a) fine-tuned on domain-specific data with cross-entropy loss (without human-saliency), and (b) fine-tuned using CYBORG loss (i.e.,with human-saliency).
Thus, answering RQ1, we conclude that MENTOR pre-training is an effective method to improve the generalization capabilities in iris PAD, synthetic face detection, and chest x-ray-based diagnosis tasks.
5.2 Answering RQ2 (fine-tuning effectiveness)
To answer RQ2, we qualitatively evaluated model’s validation accuracy during training for all combinations of the domain and network architecture. As seen in Figs. 6-8, MENTOR helps reduce training time by bringing the models to the convergence point faster. We observe higher validation accuracy and smaller standard deviations in case of MENTOR pre-training across all three model architectures. These results indicate that MENTOR provides more efficient use of training data and human annotations, both of which are costly to obtain. Thus, answering RQ2 we conclude that MENTOR increases training efficiency in three domains considered in this work.
@@ 6 Conclusion
This paper introduces MENTOR, a novel and intuitively straightforward pre-training method that embeds the “understanding” of domain-specific human salient features into the model, different from previous human perception-guided training approaches, which align the spatial correspondence between human-sourced and model-sourced saliences. To do that, we first trained an autoencoder to learn associations between human-salient features and images of both classes. Once the latent representation between the input image and the human saliency is learned, we used the encoder part to build a domain-specific classifier. We showed that using MENTOR pre-training results in performance gains in classification tasks, even compared to a model whose weights were first transferred from a network trained on a vast number of images (ImageNet). We also demonstrated that MENTOR pre-training achieved better generalization than state-of-the-art approach to incorporate human salience during training via specially-designed loss function. Finally, we show that initialization of classification models with MENTOR weights improves training efficiency by converging slightly faster than models initialized with non-human-perception-driven weights. Future work is dedicated towards exploring the minimum amount and quality (e.g.,experts- vs non-experts-sourced) of human saliency data necessary to benefit from MENTOR pre-training paradigm.

### v3

@@ 5 Results
         ResNet152 0.840$\pm$0.01 0.854$\pm$0.01 0.840$\pm$0.011 0.856$\pm$0.01
         Inception-V4 0.840$\pm$0.01 0.830$\pm$0.02 0.791$\pm$0.018 0.851$\pm$0.00
         EfficientNet-7b 0.817$\pm$0.01 0.820$\pm$0.01 0.829$\pm$0.009 0.838$\pm$0.01
Table 2: MENTOR pretraining combined with each baseline method. Means and standard deviations of AUROC scores are reported across 10 independent runs using a ResNet152 backbone.
Initialization Weights ImageNet [13] MENTOR (ours)
Iris PAD
CYBORG [4] 0.897$\pm$0.020 0.940$\pm$0.026
UNET+Gaze [22] 0.892$\pm$0.030 0.914$\pm$0.017
Synthetic Face Detection
CYBORG [4] 0.508$\pm$0.050 0.549$\pm$0.014
UNET+Gaze [22] 0.432$\pm$0.042 0.439$\pm$0.059
Chest X-ray-based diagnosis
CYBORG [4] 0.854$\pm$0.010 0.867$\pm$0.006
UNET+Gaze [22] 0.840$\pm$0.011 0.868$\pm$0.005
To answer both research questions, we compare MENTOR-trained models with those (a) pretrained using a classical learning approach, initialized with ImageNet weights, and (b) trained with a state-of-the-art human perception-guided approaches using the CYBORG loss [4] and UNET+Gaze [22].
5.1 Improvement of Generalization
MENTOR pretraining compares very favorably to models initialized with ImageNet weights, as well as to models trained with a state-of-the-art human-perception training paradigms using the same data and the same human saliency maps (see Tab. 1). MENTOR achieves improved average performance in all but one training configuration in Synthetic Face Detection (ResNet152 trained using cross-entropy, AUC=0.540), where it was marginally worse (0.539). MENTOR displayed high performance for Iris PAD across all backbones, with the best being ResNet. Inception and EfficientNet were equally strong in raising generalization performance with MENTOR (0.899 and 0.901, respectively), with CYBORG (0.794) and UNET+Gaze (0.807) performing the poor on Inception-based backbones, even lagging behind cross-entropy (0.879). MENTOR performed well in synthetic face detection, offering large improvements over existing saliency-based methods (CYBORG, UNET+Gaze), who struggled to attain above random chance. These findings suggest that existing human saliency training methods used in synthetic face detection struggle in smaller training sets (only 765 samples). Finally, MENTOR showed consistently strong performance in chest x-ray diagnosis across all models using MENTOR, with the largest gains compared to existing methods coming from Inception. Thus, answering RQ1, we conclude that MENTOR pretraining is an effective method to improve the generalization capabilities in iris PAD, synthetic face detection, and chest x-ray-based diagnosis tasks.
5.2 Combining MENTOR with other methods
To answer RQ2, we evaluated two human-guided training methods (CYBORG & UNET+Gaze) using MENTOR’s pretrained weights. As seen in Tab. 2, MENTOR greatly increases the generalization performance of both methods across all domains. MENTOR pretraining coupled with the CYBORG loss function resulted in a 0.043 increase in Iris PAD, and 0.041 increase for synthetic face detection. UNET+Gaze had a substantial benefit in the chest x-ray diagnosis task. These results indicate that MENTOR pretraining provides human perception-guided methods with salient-aware kernels that more easily adapt to a better solution space than conventional methods. In addition, these results indicate MENTOR’s strengths in complementing existing saliency-based training strategies, further contributing to its novelty and expanding its potential widespread use. Thus, answering RQ2 we conclude that MENTOR pretraining can be easily adapted to other saliency-based methods to increase the generalization performance in the three domains considered in this work.
Table 3: Ablation study with different sized ResNets for all tasks. AUC Means and standard deviations are reported across 10 independent runs, as well as the number of trainable parameters (M). The best results for each backbone are bolded, second best are underlined.
Backbone Params (M) Iris PAD Face Chest
ResNet18
Cross-entropy 11.7 0.862$\pm$0.005 0.475$\pm$0.038 0.834$\pm$0.004
CYBORG [4] 11.7 0.890$\pm$0.019 0.614$\pm$0.028 0.834$\pm$0.010
UNET+Gaze [22] 14.3 0.818$\pm$0.064 0.412$\pm$0.053 0.819$\pm$0.012
MENTOR (ours) 11.7 0.875$\pm$0.012 0.450$\pm$0.038 0.847$\pm$0.006
ResNet34
Cross-entropy 21.8 0.887$\pm$0.020 0.510$\pm$0.057 0.823$\pm$0.017
CYBORG [4] 21.8 0.894$\pm$0.017 0.559$\pm$0.055 0.841$\pm$0.007
UNET+Gaze [22] 24.4 0.789$\pm$0.097 0.399$\pm$0.049 0.815$\pm$0.008
MENTOR (ours) 21.8 0.893$\pm$0.015 0.534$\pm$0.053 0.837$\pm$0.01
ResNet50
Cross-entropy 25.6 0.874$\pm$0.017 0.500$\pm$0.055 0.833$\pm$0.011
CYBORG [4] 25.6 0.875$\pm$0.028 0.531$\pm$0.050 0.846$\pm$0.010
UNET+Gaze [22] 32.5 0.879$\pm$0.022 0.411$\pm$0.040 0.842$\pm$0.020
MENTOR (ours) 25.6 0.912$\pm$0.028 0.506$\pm$0.044 0.844$\pm$0.006
ResNet152
Cross-entropy 60.2 0.879$\pm$0.02 0.540$\pm$0.04 0.840$\pm$0.01
CYBORG [4] 60.2 0.897$\pm$0.02 0.508$\pm$0.05 0.854$\pm$0.01
UNET+Gaze [22] 67.2 0.892$\pm$0.03 0.432$\pm$0.042 0.840$\pm$0.011
MENTOR (ours) 60.2 0.912$\pm$0.02 0.538$\pm$0.05 0.856$\pm$0.01
6 Ablation Study
In this section, we provide extended experimentation and analysis of MENTOR under a variety of different benchmarks and scenarios. While these experiments are not core to our research contributions, they bolster the overall capabilities of our method. First, we benchmark MENTOR against the Teacher-Student Paradigm, which is a SOTA method used to extend the limited uses of human saliency. Second, we investigate both qualitatively and quantitatively the saliency maps offered by each of the methods. Finally, we explore MENTOR’s capabilities on a variety of smaller ResNet backbones.
6.1 MENTOR & Teacher-Student Framework
Table 4: Ablation study containing MENTOR’s performance against the Teacher-Student paradigm for the Iris PAD task. For readability purposes, the table is reproduced following Crum et al., and follows the same training configurations and evaluation measures. Baseline 1 indicates small training set accompanied by human salience, Baseline 2 indicates larger training sets without human salience. The Optimal AI Student [8] used a chain of models to provide CNN architectures with various forms of model saliency for larger training sets. MENTOR maintains competitive across half of the architecture without any additional fine-tuning. Best AUC is bolded, whereas second best AUC is underlined.
Model Baseline 1 [8] Baseline 2 [8] Optimal AI Student [8] MENTOR (ours)
DenseNet121 0.920$\pm$0.017 0.917$\pm$0.017 0.950$\pm$0.013 0.931$\pm$0.013
ResNet50 0.854$\pm$0.031 0.905$\pm$0.013 0.920$\pm$0.022 0.941$\pm$0.017
Xception 0.852$\pm$0.018 0.948$\pm$0.008 0.952$\pm$0.003 0.947$\pm$0.004
Inception-v3 0.888$\pm$0.018 0.905$\pm$0.029 0.947$\pm$0.010 0.949$\pm$0.009
Although our focus was on raising the generalization performance in limited training samples, MENTOR’s two round training strategy can be utilized to synthetically generate saliency for subsequent training samples. To explore this idea, we compared MENTOR’s salience generated from Step 1 the Teacher-Student paradigm [8]. Due to time and compute constraints, we included experiments for the Iris PAD task following the same configurations and training splits offered by [8] (see Tab. 4). Without any fine-tuning, “student” models trained using MENTOR’s saliency was able to outperform SOTA performance in half of the published architectures, with dominant gains with ResNet50 (0.920 to 0.941) and marginal improvement with Inception-V3 (0.947 to 0.947). MENTOR remained competitive for Xception-based architectures (0.952 to 0.947), but lagged behind for DenseNet (0.950 to 0.931). Though not a main focus of our work, we are able to indicate the versatility of our method.
6.2 Analysis of Saliency Maps
Figure 3: Same as Fig. 2, except UNET+Gaze [22], CYBORG [4], and Cross-entropy (Xent) saliency maps are shown. MENTOR and UNET+Gaze saliency are generated from their respective decoder, whereas Cyborg and Xent are generated using a Class Activation Mapping (CAM). Samples were generated from the validation set.
Finally, we included both a qualitative and quantitative analysis comparing the saliency generated by MENTOR and the other methods included in this paper.
Qualitative evaluation
An extended visualization of Fig. 2 can be found in Fig. 3 for more examples. As discussed previously (Sec. 3), saliency-based training is traditionally trained jointly, where the network is tasked with satisfying both the classification and human-salient (i.e. often reconstruction, or minimizing the difference between human and model saliency maps) components. However, with more strenuous tasks, the network is often forced to make a trade off between classification performance or the adherence towards human saliency (otherwise simply holistic task understanding). Occasionally, these components may altogether prove to be incompatible with existing techniques when trained jointly. Fig. 3 shows UNET+Gaze suffering from information collapse during its reconstruction of the human saliency map, and instead favoring higher classification performance. Likewise, CYBORG produces a low-resolution saliency map that misses key features found within the human annotation. Instead, MENTOR alleviates these concerns by employing a two-round training strategy, where first the task-specific representations are built into the model (resulting in high reconstruction quality), and second associations are made between these feature representations and the classification labels.
Table 5: Ablation study of measuring the Salience Entropy ($S_{entropy}$) of the Class Activation Mapping (CAM) generated from the test set [9]. Means and standard deviations are reported for a ResNet152 backbone across several different methods. Lower ($S_{entropy}$) typically suggest more compact or focused saliency.
Method Iris PAD Face Chest
Cross-entropy 0.958$\pm$0.018 0.959$\pm$0.017 0.960$\pm$0.016
CYBORG [4] 0.963$\pm$0.016 0.942$\pm$0.021 0.938$\pm$0.023
MENTOR (ours) 0.961$\pm$0.017 0.959$\pm$0.017 0.960$\pm$0.018
Quantitative evaluation
In addition, we included some quantitative statistics involving the saliency maps generated from cross-entropy, CYBORG, and MENTOR trained models offered by [9]. These measures provide a toolkit to access the explainability of model salience, especially within the context of synthetic face detection. Due to time constraints, we measured the Salience Entropy ($S_{entropy}$) of the Class Activation Mappings (CAM) generated on the test set [9, 48]. Salience Entropy attempts to measure the compactness or focus of the saliency maps. Initial observations appear to indicate that MENTOR may have slightly higher average $S_{entropy}$ than existing saliency-based methods, contrary to previous findings in synthetic face detection [9]. However, due to the sparsity of saliency (i.e. eye-tracking associated with chest x-rays, and attack-type specific annotations for iris PAD related tasks, new ways of quantitatively comparing saliency may be necessary. In this work, we used all available human saliency during training, so there is simply no way to measure the efficacy between human and model salience without collecting additional data, saving this for future work.
6.3 MENTOR on Smaller Architectures
Our core findings suggest that MENTOR can be an effective strategy to raise generalization in larger CNN models. We also included additional experimentation that support MENTOR’s ability to raise generalization in smaller models (see Tab. 3). MENTOR provides the highest gains in generalization performance across ResNet152 and ResNet50 backbones, while being competitive with the CYBORG loss for ResNet32 and ResNet18 for Iris PAD and Chest x-ray anomalies. Note that no additional changes in training configurations were made between all ResNet experiments. For future experiments, likely changes with the learning rate will be the most beneficial for using MENTOR on smaller models (otherwise using different optimizers). As shown previously with RQ2, MENTOR pairs well with strong regularization loss components (see Tab. 2). Likely a combination between MENTOR and loss function components (i.e. CYBORG) could greatly raise the generalization performance depending on the task.
@@ 7 Conclusion
This paper introduces MENTOR, a novel and intuitively straightforward pretraining method that embeds the “understanding” of domain-specific human salient features into the model, different from previous human perception-guided training approaches, which align the spatial correspondence between human-sourced and model-sourced saliences. To do that, we first trained an autoencoder to learn associations between human-salient features and images of both classes. Once the latent representation between the input image and the human saliency is learned, we used the encoder part to build a domain-specific classifier. We showed that using MENTOR pretraining results in performance gains in classification tasks, even compared to a model whose weights were first transferred from a network trained on a vast number of images (ImageNet). We also demonstrated that MENTOR pretraining achieved better generalization than state-of-the-art approach to incorporate human salience during training via specially-designed loss function. We show that other human perceptual-guided training techniques initialized with MENTOR weights improves generalization performance when compared with traditional, non-human-perception-driven weights (ImageNet). Finally, we included a detailed ablation study highlighting the versatility of our method across the Teacher-Student Paradigm, salience generation, and various model sizes.
Acknowledgments
This work was supported by the U.S. Department of Defense (Contract No. W52P1J-20-9-3009) and by the National Science Foundation (Grant No. 2237880). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation, the U.S. Department of Defense or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes, notwithstanding any copyright notation here on. We would also like to acknowledge and extend our thanks to Jacob Piland for his work in curating and preparing the chest X-ray dataset featured in this paper.